OpenStudy (anonymous):

Find -3A+6B A=[-3 5 -6] B=[-2 6 7] [9 -5 3] [2 -1 -6]

OpenStudy (unklerhaukus):

\[\mathbf A=\begin{bmatrix}-3&5&-6\\9&-5&3\end{bmatrix}\] \[\mathbf B=\begin{bmatrix}-2&6&7\\2&-1&-6\end{bmatrix}\] \[-3\mathbf A=\begin{bmatrix}-3\times-3&-3\times5&-3\times-6\\-3\times9&-3\times-5&-3\times3\end{bmatrix}\] \[6\mathbf B=\begin{bmatrix}6\times-2&6\times6&6\times7\\6\times2&6\times-1&6\times-6\end{bmatrix}\]

OpenStudy (anonymous):

@UnkleRhaukus shouldn't all of the 3's be negative?

OpenStudy (unklerhaukus):

yeah i fixed that now*

OpenStudy (anonymous):

@UnkleRhaukus oh ok. so do i just multiply them then subtract?

OpenStudy (unklerhaukus):

yeah, the element on the top left will be (−3×−3)+(6×-2)

OpenStudy (anonymous):

@UnkleRhaukus 15x-2?

OpenStudy (unklerhaukus):

how did you get that?

OpenStudy (anonymous):

no 3x-5..?

OpenStudy (anonymous):

@UnkleRhaukus i'm so confused.

OpenStudy (unklerhaukus):

ok, lets go slower what does this simplify to become? \[-3\mathbf A=\begin{bmatrix}-3\times-3&-3\times5&-3\times-6\\-3\times9&-3\times-5&-3\times3\end{bmatrix}\]

OpenStudy (anonymous):

9 -15 18 -27 15 -9 @UnkleRhaukus

OpenStudy (unklerhaukus):

good

OpenStudy (unklerhaukus):

and what about \[6\mathbf B=\begin{bmatrix}6\times-2&6\times6&6\times7\\6\times2&6\times-1&6\times-6\end{bmatrix}\]

OpenStudy (anonymous):

-12 36 42 12 -6 -36 @UnkleRhaukus

OpenStudy (anonymous):

i got -3 21 60 -15 9 -45 as my answer @UnkleRhaukus

OpenStudy (unklerhaukus):

right so you have -3A + 6B = [ 9 -15 18 ] [ -12 36 42 ] [ -27 15 -9 ] + [ 12 -6 -36 ] = [ 9-12 -15+36 18+42 ] [-27+12 15-6 -9+-36 ]

OpenStudy (unklerhaukus):

that's right!

OpenStudy (anonymous):

thanks for your help :) @UnkleRhaukus